A wheel rotates with a constant acceleration of 2.0 rad/s^2. If the wheel starts from rest , how many revolutions will it make in the first 10 seconds?

Grade:11

1 Answers

Parul singh
7 years ago
Angular acceleration= omega/time Omega= 2×10=20 theta=( final- initial angular velocity/2)×t =100 no of revolution=theta /2π =15.9